OpenStudy (anonymous):

Solve for x and check. Enter only a number. 3^2x = 27

OpenStudy (anonymous):

3^2x=3^3 2x=3 x=3/2

OpenStudy (ja1):

^

OpenStudy (anonymous):

not the answer @marsss

OpenStudy (anonymous):

I think 9x=27 x=3.....but im not suree